OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5).

OpenStudy (anonymous):

g(-5) = -2 (-5) +4 = 10 + 4 = 14

OpenStudy (saifoo.khan):

g(x) = −2x + 4 g(-5) = −2(-5) + 4 g(-5) = 10 + 4 14.